At the end of September, a barrel of light crude sold for almost $70 compared to a price near $30 a barrel in January of 2004. To answer the following questions. Assume that bond trade s expect inflation to rise from 3 percent in 2005 (history) to 5 perc in both 2006 and 2007 (expected inflation). Also traders expect the Zambian economy to enter a recession in 2007. Assume the prior to the recent run up in oil prices, bond traders had expected inflation to remain stable in 2006-2007 at 3 percent.

I) using a model of supply and demand for one year T-bills, illustrate and explain the situation impact of a recession (a business cycle contraction) if bond traders expect it that this recession will occur in 2007, what do you expect to happen to yields on one year T-bills in 2007?
1
Expert's answer
2021-09-13T11:21:14-0400

A Treasury bill, often known as a T-bill, is a short-term government debt security that matures in less than a year. Treasury bills, unlike many other financial products, do not pay interest to investors on a regular basis. Rather, the banknotes are sold at a lower price than they would be if they were redeemed.

A recession is defined as a major drop in economic activity across the economy that lasts more than a few months and is reflected in real GDP, real income, employment, industrial output, and wholesale-retail sales. A recession starts when the economy reaches its peak of activity and concludes when it hits its lowest point. The economy is expanding between the trough and the high. 

Treasuries must also contend with inflation, which is a measure of the rate at which prices rise in the economy. Even though T-Bills are the most liquid and stable debt asset available, investors are less likely to purchase them when inflation is higher than the T-bill yield. As a result, during inflationary periods, T-bill values tend to decline as investors sell them in favor of higher-yielding assets.


Justification

To find out how much a Treasury bill is worth, we must compare its par value to its face value. We'll also need to utilize the maturity term to convert the return to an annual percentage because investment returns are most helpful when presented on an annualized basis. So, the first step is to acquire some data: you'll need to know the Treasury bill's purchase price, purchase date, and maturity date. To annualize the return, you'll need to calculate the number of days to maturity based on this information. 

Suppose 

If a Treasury bill at a price of 99

expected inflation in 2007 is 5%

Now we will calculate the yield on one year T- Bills in 2007

The first step is to subtract the price of the T-bill from 100 and divide the result by the price. This figure tells the yield on a T-bill over its entire maturity period. To convert this amount to a percentage, multiply it by 100.

"Yield=\frac{100−price}{price}×100\\
Yield=\frac{100−99}{99}×100\\
Yield=\frac{1}{99}×100\\
Yield=1.01%"